# Eaman Jahani

### v1

@@ B.5.2 Prompt Length
We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt. In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9952$, 95% CI: [0.99507, 0.99543]).
@@ B.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
B.5.6 Inversions of Successive Measures
[…]
B.5.7 Prompt Composition
@@ B.5.8 Strategic Shifts
It is worth noting that Torricelli et al. (2023) uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift. We instead use the task-level mean because in our setting, it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
@@ C.2 Z-Scoring
We find statistically significant evidence for differences in the variability of performance across the 15 target images used in our experiments, which we discuss in Section F. In the main text, we also showed that performance increases across the attempts. To test whether our results are in some way due to this image-level or attempt-level variations, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment. Formally, this is:
@@
where $\textrm{CosineSim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$. The mean and standard deviation are computed per each image-attempt, but over the DALL-E 2 and DALL-E 3 treatment arms. We also use the rescaling above to test the robustness of our DreamSim-based analyses.
@@ C.3 Accounting for Model Stochasticity
The output that generative AI models return in response to a given prompt is stochastic. The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E 3 API at the time of our experiment. To account for this model stochasticity, we generated 10 images for each prompt submitted by participants for all arms. We were then able to calculate the similarity between each replication and its corresponding target image, and calculate an “expected” CLIP cosine similarity and DreamSim score for each prompt by averaging over these samples.
We generated these additional samples for both the original prompts on their assigned treatment arms, as well as replaying on the counterfactual arms, as introduced in Figure 2 in the main text. Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling. As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts. This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the under the ”replaying” procedure (Section B.B.3.3).
@@ D.1 Task Performance and ATEs
Next, the bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane. [^5] In Section D, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified. To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors:
@@
  • $\hat{\beta_{1}}=0.0015$, $\hat{SE}(\beta_{1})=0.0003$, $p=9.678\times 10^{-7}$
  • $\hat{\beta_{2}}=0.0126$, $\hat{SE}(\beta_{2})=0.003$, $p=1.7\times 10^{-4}$
  • $\hat{\beta_{3}}=0.00086$, $\hat{SE}(\beta_{3})=0.0004$, $p=0.0418$
@@ D.2 Prompt Characteristics
Figure 1C compares the prompt length and prompt similarity of the two models. To generate these results, we first remove any prompt that does not resemble a good-faith attempt according to the sample construction procedure detailed in Section B.B.4. The prompt length is the average number of words per model and iteration stratified by the reference image. The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section B.B.5.3 for details on similarity calculations). The color scale in figure 1C shows the stratified average similarity to the previous prompt across all users per each model. We find that superior model users write prompts that, on average, have $\beta=0.0166$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.02$).
Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from the superior model participants were, on average, more similar than the prompts of the inferior model participants. For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section B.B.5.3, as the dependent variable. When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=-0.0169$ smaller than inferior users ($p=0.008$). Standard errors are cluster-robust at the participant level.
@@ E.1.1 Overall ATEs
In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.261$ standard deviations (95% CI = [0.180, 0.342]) closer to the target image ($\Delta DreamSim=0.0304$, $p<10^{-9}$) than those produced by participants using DALL-E 2 (the inferior model).
@@ E.1.2 Figure 1
  • $\hat{\beta_{1}}=0.0039$, $\hat{SE}(\beta_{1})=0.00048$, $p=4.3\times 10^{-16}$
  • $\hat{\beta_{2}}=0.0179$, $\hat{SE}(\beta_{2})=0.0060$, $p=0.0029$
  • $\hat{\beta_{3}}=0.0023$, $\hat{SE}(\beta_{3})=0.00071$, $p=0.001$
@@ E.1.3 Figure 2
Decomposing the ATE as measured in terms of DreamSim, we find similar results to those in the main text. The model effect accounts for 55% of the ATE ($\Delta DreamSim=0.0168$, $p<10^{-10}$), whereas the prompting effect accounts for 45% of the ATE ($\Delta CoSim=0.01364$, $p=0.0053$).
@@ E.2 Z-score-based Analysis
As we discuss in Section B, we repeat all main-text analyses with the within-image Z score of CLIP-cosine similarity to better account for variation between images. The results hold across the board and sometimes, the differences between the superior and inferior model are even starker than when using cosine similarity.
@@ E.2.1 Overall ATEs
As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.20$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.20, 95% CI = [0.126, 0.284]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model). Standard errors are cluster robust by participants.
@@ E.2.2 Figure 1
On average, participants using the superior model produced images that were $z=0.20$ standard deviations closer (the ATE) to the target image than those using the inferior model. Like in the main text with CLIP cosine similarity, this treatment effect increased as participants made successive attempts to replicate the target image.
@@
  • $\hat{\beta_{1}}=0.02012$, $\hat{SE}(\beta_{1})=0.0041$, $p=1.19\times 10^{-6}$
  • $\hat{\beta_{2}}=0.1439$, $\hat{SE}(\beta_{2})=0.0466$, $p=0.002$
  • $\hat{\beta_{3}}=0.01253$, $\hat{SE}(\beta_{3})=0.0058$, $p=0.03$
@@ E.2.3 Figure 2
When we decompose the ATE into the model effect ($z=0.0895$; $p=1.02\times 10^{-5}$) and prompting effect ($z=0.1139$; $p=0.0044$), they account for 44% and 56% of the treatment effect, respectively. And when we replay the inferior model prompts on the superior model, the difference in similarity to the target from these prompts played on the inferior is not statistically significant and close together ($z=-0.03$; $p=0.43$). In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z score of the cosine similarity.
@@ Appendix F Pre-registration
Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses. In our pre-registration, we declared the intent to conduct each of our analyses using eight possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in all of four ways:
@@
  • GPT-4V baseline rescaling: We rescale the outcome variable by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image.
We intend to complete each pre-registered analysis with each of these eight possible outcome variables. Although we have not yet done so, we have no reason to believe that our results will not be robust to these different transformations of our data. However, we did not want to delay the publication of our preprint until all of these analyses were complete. Thus, we are posting preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes. As we complete more of these analyses, we will update this supplement. We also intend to post full results, data, and replication code online shortly.
@@ F.1 Hypotheses and Results
    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, writing frequency, and self-reported occupational skills of critical thinking and social perceptiveness.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender, age, and the critical thinking occupational skill.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    No rescaling: We find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing. Token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were positively associated with performance, while embedding variance and the number of topical transitions negatively correlated with performance. Taken together, these findings suggest that more successful users engaged in more exploitation and wrote prompts that were similar to one another. We also found that longer prompts were associated with higher performance, as shown in the main text.
    We also find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance. In contrast, in cases where previous performance was high, higher exploitation monotonically increased performance. Using token sort ratio with the previous prompt as the measure of exploration also generated similar results. However, when using binary measures of exploration, i.e. topical transition or containing the previous prompt, we found that exploitation is associated with higher performance in the next iteration regardless of the bracket. In particular, we did not find the non-linear pattern we observed with continuous measures of exploration in the low-performing group, likely due to the binary nature of this measure that fails to distinguish between low, medium, and high levels of exploration/exploitation.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    No rescaling:Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
To test this hypothesis, we estimated a linear model per each demographic trait as the independent variable, treatment arm fixed effects, and the prompting behaviors outlined below and described in-depth in Section B.B.5 as the outcome variables of interest. To account for multiple testing, we adjusted the p-values for all these models according to Benjamini-Hochberg procedure.
@@
  • Strategic Shifts: We observed statistically significant differences based on age and reported programming frequency. Older participants and those with high programming frequency demonstrated increased topical transitions across their prompts.
  • Successive Prompt Token Sort Ratio: We found significant differences by age, programming frequency, and writing frequency. Older users, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
  • Successive similarity: We found significant differences by age, gender, generative AI outlook, programming skill/frequency and writing frequency. On average, older users, males, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, and those who write a lot, those who self-describe as being good at writing instructions and writing with a lot of clear imagery write prompts that are less similar to each other (as measured by cosine similarity) on average.
  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and writing frequency. Older users and those who write frequently are less likely to write prompts that exactly contain the previous. In contrast, males and those with a positive outlook on generative AI are more likely to keep using their previous prompt.
@@
    No rescaling: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt when conducting two-sample t-tests stratified by treatment arm. When we test differences in the trichotomous variable, the difference between the top bracket and the other two brackets is significant, but we do not find significant differences between the bottom and middle brackets. However, we do not find a statistically significant relationship between the observed z-score of the prompt and performance in the next iteration when we estimate a linear model. This may be due to the non-linearity of the relationship between these variables. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average subsequent performance.
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    No rescaling: Forthcoming
    Z-score rescaling: Forthcoming
    Percentile rank rescaling: Forthcoming
    GPT-4V baseline rescaling: Forthcoming
@@
    Prompt Length: Forthcoming
    Successive Similarity: We find statistically significant evidence that increases in the z-score of the image realized in a previous prompt caused increases in successive similarity. When we compare the lowest z-score bracket (worse than expected images) to the highest z-score bracket (better than expected images), we observe a statistically significant difference as well. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. Additionally, when we estimate the effect at the user-level, we do not find statistically significant effects of the z-score of the first prompt on average successive similarity.

### v2

@@ B.5.2 Prompt Length
We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt. In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9954$, 95% CI: [0.99528, 0.99560]).
@@ B.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
B.5.6 Prompt Composition
@@ B.5.7 Strategic Shifts
It is worth noting that Torricelli et al. (2023) uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift. We instead use the task-level mean because in our setting, as it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
@@ C.2 Z-Scoring
We find statistically significant evidence for differences in the variability of performance across the 15 target images used in our experiments, which we discuss in Section F. In the main text, we also showed that performance increases across the attempts. To test whether our results are in some way due to this image-level or attempt-level variation, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment. Formally, this is:
@@
where $\textrm{CosineSim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$. The mean and standard deviation are computed per each image-attempt, but over the DALL-E 2 and DALL-E 3 treatment arms. We also use the rescaling above to test the robustness of our DreamSim-based analyses. Almost all the robustness analysis reported here use the Z-score scaled measure of performance within each image-attempt set. The only exception is any analysis that examines improvements and prompts across attempts, e.g. figure 1B, where Z-scores are computed only within each target image, and across all attempts of that image.
@@ C.3 Accounting for Model Stochasticity
The output that generative AI models return in response to a given prompt is stochastic. The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E API at the time of our experiment. To account for this model stochasticity, we generated 10 images for each prompt submitted by participants for all arms. We were then able to calculate the similarity between each replication and its corresponding target image, and calculate an “expected” CLIP cosine similarity and DreamSim score for each prompt by averaging over these samples. Given the replicated images, we can also calculate the standard deviation of cosine similarity induced by this stochasticity. With the expected cosine similarity and its standard deviation per prompt, we can compute a normalized Z-score for the observed image relative to its replication distribution. This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and will be used for further analysis in section F.
We generated these additional samples for both the original prompts on their assigned treatment arms, as well as replaying on the counterfactual arms, as introduced in Figure 2 in the main text. Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling. As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts. This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section B.B.3.3).
@@ D.1 Task Performance and ATEs
Next, the bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane. [^5] In Section D, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified. To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
@@
  • $\hat{\beta_{1}}=0.0011$, $\hat{SE}(\beta_{1})=0.0003$, $p=0.0004$
  • $\hat{\beta_{2}}=0.0120$, $\hat{SE}(\beta_{2})=0.0037$, $p=0.0013$
  • $\hat{\beta_{3}}=0.0010$, $\hat{SE}(\beta_{3})=0.0004$, $p=0.0227$
@@ D.2 Prompt Characteristics
Figure 1C compares the prompt length and prompt similarity of the two models. To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section B.B.4. The prompt length is the average number of words per model and iteration stratified by the reference image. The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section B.B.5.3 for details on similarity calculations). The color scale in figure 1C shows the stratified average similarity to the previous prompt across all users per each model. We find that superior model users write prompts that, on average, have $\beta=0.0184$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.0236$).
Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the inferior model participants. For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section B.B.5.3, as the dependent variable. When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$). Standard errors are cluster-robust at the participant level.
@@ E.1.1 Overall ATEs
In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.238$ standard deviations (95% CI = [0.152, 0.324]) closer to the target image ($\Delta DreamSim=0.0306$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model).
@@ E.1.2 Figure 1
  • $\hat{\beta_{1}}=0.0034$, $\hat{SE}(\beta_{1})=0.0005$, $p=1.3\times 10^{-12}$
  • $\hat{\beta_{2}}=0.0200$, $\hat{SE}(\beta_{2})=0.0061$, $p=0.0011$
  • $\hat{\beta_{3}}=0.0024$, $\hat{SE}(\beta_{3})=0.0007$, $p=0.0009$
@@ E.1.3 Figure 2
Decomposing the ATE as measured in terms of DreamSim, we find similar results to those in the main text. The model effect accounts for 54.4% of the ATE ($\Delta DreamSim=0.0166$, $p<10^{-7}$), whereas the prompting effect accounts for 45.4% of the ATE ($\Delta DreamSim=0.01390$, $p=0.014$).
@@ E.2 Z-score-based Analysis
As we discuss in Section C.2, we repeat all main-text analyses with the within-image-attempt Z-score of CLIP-cosine similarity to better account for variation between images and attempts. When comparing across attempts, the Z-score is computed within each image, as mentioned in section C.2. The results hold across the board and sometimes, the differences between the superior and inferior model are even starker than when using cosine similarity.
@@ E.2.1 Overall ATEs
As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.19$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.19, 95% CI = [0.100, 0.271]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2 (the inferior model). Standard errors are clustered at the participant level.
@@ E.2.2 Figure 1
On average, participants using the superior model produced images that were $z=0.19$ standard deviations closer (the ATE) to the target image than those using the inferior model. Like in the main text with CLIP cosine similarity, this treatment effect increased as participants made successive attempts to replicate the target image.
@@
  • $\hat{\beta_{1}}=0.0129$, $\hat{SE}(\beta_{1})=0.0038$, $p=0.0007$
  • $\hat{\beta_{2}}=0.1250$, $\hat{SE}(\beta_{2})=0.0457$, $p=0.0064$
  • $\hat{\beta_{3}}=0.0128$, $\hat{SE}(\beta_{3})=0.0053$, $p=0.015$
@@ E.2.3 Figure 2
When we decompose the ATE into the model effect ($z=0.0791$; $p=8.35\times 10^{-6}$) and prompting effect ($z=0.1046$; $p=0.016$), they account for 43% and 56% of the treatment effect, respectively. And when we replay the inferior model prompts on the superior model, the difference in similarity to the target from these prompts played on the inferior is not statistically significant and close together ($z=-0.033$; $p=0.45$). In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z-score of the cosine similarity.
@@ Appendix F Pre-registration
Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses. In our pre-registration, we declared the intent to conduct each of our analyses using six777see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables. possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
@@
We intend to complete each pre-registered analysis with each of these eight possible outcome variables. Although we have not yet done so, we have no reason to believe that our results will not be robust to these different transformations of our data. However, we did not want to delay the publication of our preprint until all of these analyses were complete. Thus, we are posting a preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes. As we complete more of these analyses, we will update this supplement. We also intend to post full results, data, and replication code online shortly.
@@ F.1 Hypotheses and Results
    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, active listening and quality control.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as job skills, and report little usage of generative AI or imagery writing skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with self-reported occupational skill of quality control, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
    Z-score rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, quality control, technology design, social perceptiveness and troubleshooting.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and troubleshooting as job skills, and report little usage of generative AI, little imagery writing skill, some programming skill, and some instruction writing skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with a graduate degree, those with self-reported occupational skill of quality control and technology design, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
    Percentile rank rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, and self-reported occupational skill of critical thinking.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as a job skill, and report little usage of generative AI, and some programming skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
@@
    No rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skill, and self-reported occupational skills of critical thinking, learning strategies, technology design and quality control.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies occupational skill.
    Z-score rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skills, education, and self-reported occupational skills of critical thinking, social perceptiveness, learning strategies, technology design and quality control.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, those with graduate degrees, and those with technology design, quality control and learning strategies as occupational skills performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
    Percentile rank rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery writing skill, education, and self-reported occupational skills of social perceptiveness, learning strategies, and technology design.
    When we fit a linear model to evaluate directionality, we find that those who use social perceptiveness as a job skill, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, and those with technology design and learning strategies as occupational skills performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, generative AI use, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
@@
    No rescaling: At the user-level, we find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing. Token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were positively associated with performance, while embedding variance and the number of topical transitions negatively correlated with performance. Taken together, these findings suggest that more successful users engaged in more exploitation and wrote prompts that were similar to one another. We also found that longer prompts were associated with higher performance, as shown in the main text.
    At the user-prompt level, we find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance. In contrast, in cases where previous performance was high, higher exploitation monotonically increased performance. Using token sort ratio with the previous prompt as the measure of exploration also generated similar results. However, when using binary measures of exploration, i.e. topical transition or containing the previous prompt, we found that exploitation is associated with higher performance in the next iteration regardless of the bracket. As these are binary measures of exploration, we could not replicate the non-linearity we observed with continuous measures of exploration in the low-performing group. However, we do find that topical transitions (more exploration) leads to an overall drop in performance and the drop becomes larger for higher performance in the previous attempt. We find a similar pattern when measuring exploitation by whether a prompt contains the previous one. Performance is overall higher if the prompt includes the previous one and the improvement is larger as the previous performance gets higher.
    Z-score rescaling: We find the same results as those described for the unscaled cosine similarity above. The main difference is that the non-linearity between performance and continuous measures of exploitation, i.e. TSR ratio and cosine similarity with previous prompt, observed at the user-prompt level becomes starker for the bottom two brackets of previous performance. For low-performing prompts in the previous attempt, the optimal exploitation level is in the middle and the performance greatly deteriorates for high and low levels of exploitation.
    Percentile rank rescaling: We find the same results as described above, however the non-linearity in exploration/exploitation vs performance in the bottom two brackets of previous performance is now even stronger than those observed with Z-score rescaled cosine similarity as the performance measure.
@@
    No rescaling: We find exactly the same results as those described for the unscaled cosine similarity above.
    Z-score rescaling: We find the same results as those described for the Z-score cosine similarity above. In particular, we again find that the non-linearity between performance and continuous measures of exploitation observed at the user-prompt level is stark for the bottom two brackets of previous performance.
    Percentile rank rescaling: We find the same results as described above for cosine similarity percentile rank rescaling.
@@
To test this hypothesis, we estimated a linear model per each demographic trait as the independent variable, treatment arm fixed effects, and the prompting behaviors outlined below and described in-depth in Section B.B.5 as the outcome variables of interest. To account for multiple testing, we adjusted the p-values for all these models according to the Benjamini-Hochberg procedure.
@@
  • Strategic Shifts: We observed statistically significant differences based on age, reported programming and instructional writing frequencies. Younger participants, those with low programming frequency and those with some instructional writing frequency demonstrated decreased topical transitions across their prompts.
  • Successive Prompt Token Sort Ratio: We found significant differences by age, education, programming frequency, and imagery/instructional writing frequencies. Older users, those with post-graduate degrees, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
  • Successive similarity: We found significant differences by age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency and some occupational skills. On average, older users, males, those with post-graduate degrees, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, those who report frequently writing instructions or imagery, and those with self-reported occupational skills of critical thinking and social perceptiveness write prompts that are less similar to each other (as measured by cosine similarity) on average.
  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and imagery writing frequency. Older users and those with neutral outlook toward generative AI are less likely to write prompts that exactly contain the previous. In contrast, males and those with some imagery writing skills are more likely to keep using their previous prompt.
@@
    No rescaling: We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt. When we test differences in the trichotomous variable, the difference between the top bracket and the bottom bracket in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than the bottom bracket. But we do not find significant differences between the bottom and middle brackets and between the top and middle brackets. When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration. This is due to the non-linearity of the relationship between these variables that exists separately in the negative and positive range of realized Z-score. Additionally, when we estimate the effect at the user-level, we find differences between the top and middle brackets of the first prompt realized Z-score. However, these differences are barely significant ($p=0.048$) and do not account for multiple testing. When fitting a linear model at the user-level, we do not find a a statistically significant relationship between the observed Z-score of the first prompt and average performance in the subsequent attempts.
    Z-score rescaling: The results are very similar to those described for the unscaled cosine similarity above. In particular, we find a statistically significant difference between the top bracket and the bottom bracket, with the top bracket having a higher performance in the next attempt than the bottom bracket. We don’t find statistically significant effects when comparing the middle with top or bottom brackets, or when fitting a linear model. In contrast to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when comparing the brackets on the realized Z-score of the first image nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
    Percentile rank rescaling: We find the exact same results as those explained above for Z-score cosine similarity.
@@
    No rescaling: We find similar evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt. The difference between the top bracket and the bottom/middle brackets in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than both the bottom and middle brackets. But we do not find significant differences between the bottom and middle brackets. When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration, likely due to the non-linearity that we also observed between performance measured as cosine similarity and realized Z-score of the previous prompt. When we estimate the effect at the user-level, we do not find any statistically significant relationship between realized quality of the first prompt and performance in subsequent attempts, either using the trichotomous variable on Z-score or using a linear model.
    Z-score rescaling: The results are very similar to those described for the unscaled dreamsim score above. In particular, we find a statistically significant difference between the top bracket and both the bottom and middle brackets, with the top bracket having a higher performance in the next attempt. We don’t find statistically significant effects when comparing the middle with the bottom bracket, or when fitting a linear model. Similar to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when using the trichotomous variable nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
    Percentile rank rescaling: We find the exact same results as those explained above for Z-score dreamsim similarity, with the only difference that at the user-attempt level, we observe a statistically significant difference only between the top and bottom brackets. The top bracket in realized quality no longer has a statistically significant improvement in the performance of the next attempt when compared to the middle bracket.